# Dynamic Neural Potential Field: Online Trajectory Optimization in the Presence of Moving Obstacles

### v2

@@ Abstract
Generalist robot policies must operate safely and reliably in everyday human environments such as homes, offices, and warehouses, where people and objects move unpredictably. We present Dynamic Neural Potential Field (NPField-GPT), a learning-enhanced model predictive control (MPC) framework that couples classical optimization with a Transformer-based predictor of footprint-aware repulsive potentials. Given an occupancy sub-map, robot footprint, and optional dynamic-obstacle cues, our autoregressive NPField-GPT head forecasts a horizon of differentiable potentials that are injected into a sequential quadratic MPC program via L4CasADi, yielding real-time, constraint-aware trajectory optimization. We additionally study two baselines: (NPField-D1) static-frame decomposition and (NPField-D2) parallel MLP heads for all steps.
@@
In dynamic indoor scenarios from BenchMR and on a Husky UGV in office corridors, NPField-GPT produces safer, more conservative trajectories under motion changes, while D1/D2 offer lower latency. We also compare with the CIAO* and MPPI baselines. Across methods, the Transformer+MPC synergy preserves the transparency and stability of model-based planning while learning only the part that benefits from data: spatiotemporal collision risk. Code and trained models are available at https://github.com/CognitiveAISystems/Dynamic-Neural-Potential-Field.
Keywords: Machine Learning for Robot Control, Collision Avoidance, Nonholonomic Motion Planning
1 Introduction
@@
Mobile robots that act autonomously within human-oriented environments can become significant assistants for humans. They can be applied in offices, shops, homes [1], and medical facilities. Autonomous operation requires methods for planning motion within the environment. Trajectory planning is often executed in two stages: first, a rough global path is generated via search-based [2, 3] or sampling-based [4, 5] methods; second, the global path is turned into a local trajectory under kinodynamic constraints and obstacle avoidance. The second stage is often executed online with a receding horizon strategy. Solutions for receding-horizon trajectory planning may be obtained with model predictive path integral (MPPI) [6, 7] or numerical model predictive control (MPC) [8, 9, 10, 11, 12, 13, 14, 15]. MPPI can work with arbitrary obstacle maps, but it can produce invalid or chattering solutions. Numerical MPC provides fast and stable solutions; however, it requires an analytical representation of collision danger either as a set of constraints [10] or as a cost term (repulsive potential) [^1] The term ”repulsive potential” was introduced for the artificial potential field global planner[16]. In MPC, the trajectory is biased toward safer regions according to the repulsive potential.. In this work, we consider local planning with numerical MPC in the presence of dynamic obstacles (Fig. 1).
@@
Figure 1: Overview of NPField-GPT integrated with MPC for dynamic obstacle avoidance. Left: Husky UGV operating in an office corridor. Middle: task window with occupancy grid, robot footprint, and a moving obstacle. Right: inputs—obstacle map, footprint, and dynamic-obstacle state $\left(x_{dyn},y_{dyn},\theta_{dyn}\right)$—are encoded and fed to NPField-GPT to predict a footprint-aware repulsive potential $J_{o}$. The potential is consumed inside the optimization loop by the MPC solver to compute $\mathbf{x}_{opt}[k\ldots k{+}m]$ and $\mathbf{u}_{opt}[k\ldots k{+}m]$, which are executed on the robot.
@@
Avoiding collisions with dynamic obstacles (e.g., walking human beings in an office environment) is especially challenging. It means that the obstacle map for each frame within the prediction horizon will be different. The prediction of obstacle flow is the specific task [17, 18, 19], which is out of our scope. It may be solved with various methods. Here, we consider dynamic obstacles to be predictable. The dynamic appearance of the obstacle model enlarges the computational complexity of the MPC problem due to a higher number of obstacle parameters.
We argue for a hybrid approach that keeps the strengths of both worlds: learn only the spatiotemporal collision risk, and plan with MPC. Concretely, we extend NPField to dynamic scenes by training a lightweight map/footprint encoder together with a sequence model that predicts a horizon of footprint-aware repulsive potentials. Our primary model (NPField-GPT) is an autoregressive Transformer (GPT) that conditions each step on prior steps, enabling temporally consistent, conservative predictions under motion changes and occlusions. These differentiable potentials are injected into a sequential quandratic program (SQP) for Model Predictive Control (MPC) via L4CasADi, preserving hard-constraint handling and interpretability while benefiting from data-driven risk estimation (Fig. 1).
To study accuracy–latency trade-offs important for deployment, we also consider: (NPField-D1) treating a dynamic scene as a sequence of static frames evaluated step-wise; and (NPField-D2) predicting all steps in parallel with lightweight MLP heads conditioned on dynamic-obstacle cues. Across these variants, we emphasize four workshop-aligned dimensions: Data (learning from occupancy grids plus optional obstacle states), Models (Transformer sequence prediction coupled to MPC), Evaluation (BenchMR scenarios and real Husky trials), and Safety (footprint-aware costs inside a deterministic optimizer).
Our contributions are:
  • A Transformer+MPC architecture that forecasts horizon potentials autoregressively and integrates them into real-time, footprint-aware trajectory optimization.
  • Two complementary baselines (D1/D2) enabling ablations on temporal coupling versus latency.
  • An empirical study on dynamic indoor navigation (BenchMR and Husky corridors) showing that NPField-GPT yields safer behavior while D1/D2 minimizes latency, all within real-time budgets.
  • Open-source implementation with L4CasADi integration to facilitate reproducible evaluation of learning-in-the-loop planners in everyday environments.
2 Related works
@@
2.1 Receding horizon planning and dynamic obstacles
@@
The task of local planning is to turn a fixed part of the rough global plan into a smooth trajectory segment under consideration of obstacles and kinodynamic constraints. Model Predictive Path Integral (MPPI) [6] achieves this via sampling random trajectory segments and generating a good solution based on these samples. Collision check for sample trajectories may be systematically extended to the case of dynamic obstacles [27, 28]. Computational heaviness and nondeterministic nature are disadvantages of MPPI; therefore, it may be used in cases when the process model is too complicated for MPC. Numerical MPC solves the local planning task as an optimal control problem. Obstacle avoidance is formalized as a set of constraints (e.g. [10]) or as an additional cost term of the optimization problem. This formalization leads to additional problem parameters describing obstacle properties. The number of parameters should be small to preserve MPC performance. Many approaches approximate the obstacle map with a low-dimensional model. Either obstacles or free space may be approximated with a set of simple geometric figures: points [9], circles [10, 14], rectangles [11], polylines [29], or polygons [8, 30, 15, 31]. Some of these works explicitly consider dynamic environments [29, 30]; others can be adapted straightforwardly (e.g., CIAO* [10, 11] uses independent approximations of the free space at each timestep; these approximations can reflect dynamic obstacles). The limitations of geometric approximations are computational challenges and loss of fidelity. Many works (e.g. [8, 15]) do not consider how to obtain a geometric approximation from an arbitrary obstacle map. In practice, maps are often provided as occupancy grids (a matrix projected onto the map; zero values correspond to free-space cells, while ones correspond to occupied cells). High-resolution grids have too many parameters to be passed to the MPC solver in raw form. In the next subsection, we consider approaches where obstacle data are modeled with neural networks.
2.2 Trajectory optimization with neural collision model
@@
Learning collision model for numeric trajectory optimization was considered in several works [32, 33, 34, 35, 36]. It is challenging to balance the precision and computational complexity of the collision model; therefore, existing works are constrained in terms of computation time and complexity of the maps. [32, 33, 34, 35] use neural models inspired by Neural Radiance Field [37]. These models learn the structure of a single obstacle map (in 2D or 3D) and may be used for navigation within the learned map. [32, 33] exploit network inference for trajectory optimization, while [34, 35] optimize trajectory within the learning procedure. Instead of learning a single obstacle map, the models from [36] take an obstacle representation as input. Both architectures include two submodels: the first reduces the dimensionality of the obstacle representation, while the second calculates the collision score. Therefore, the first submodel provides a compact vector of parameters for a real-time MPC solver, while the second is integrated into this solver using the L4CasADi [33] framework. The difference is that NPField utilizes occupancy grids for footprint-aware collision avoidance of a wheeled mobile robot, while [36] utilizes depth images for 3D collision avoidance of an aerial robot. To our knowledge, there are no neural obstacle models that provide MPC avoidance from dynamic obstacles in real time. [35] exploits a NeRF learning procedure for trajectory optimization in the presence of moving obstacles; however, the computational procedure takes tens of seconds, which is non-real-time. This work aims to provide a model that allows online MPC collision avoidance with a neural model of static and dynamic obstacles.
Positioning and novelty summary.
In summary, classical geometric MPC methods (e.g., CIAO* [10, 11]) require handcrafted low-dimensional approximations and do not naturally encode footprint-aware costs from raw grids. Sampling-based methods (MPPI [6, 7]) handle arbitrary maps but can be unstable and slow. Prior neural approaches either learn a single scene [32, 34, 35] or use non-map inputs [36], and most are not designed for real-time optimization in the loop. Our contribution is a family of neural potential field models that: (i) accept occupancy grids and footprint as inputs, (ii) produce differentiable repulsive potentials amenable to real-time MPC via L4CasADi, and (iii) extend to dynamic obstacles through three variants (D1, D2, NPField-GPT) with different accuracy/latency trade-offs.
3 Background
@@
Following NPField, we define the statement and notations for model predictive local planning. We consider a non-holonomic wheeled robot with a differential drive. System state vector $\mathbf{x}=\{x,y,\theta,v\}$ include 2D robot coordinates $x$ and $y$, its orientation $\theta$ and linear velocity $v$ (directed according to $\theta$). Control vector $\mathbf{u}=\{a,\omega\}$ include robot acceleration $a$ (directed according to $\theta$) and angular velocity $\omega$. We consider the model with continuous time dynamics and discrete time control (i.e., $u$ is considered to be constant within the single timestep). A formal statement for the trajectory optimization problem is the following:
$\{\mathbf{x}_{opt}[i],\mathbf{u}_{opt}[i]\}_{i=k}^{k+m}=arg\min\sum_{i=k}^{k+m}(\lVert{\mathbf{x}[i]-\mathbf{x}_{r}[i]\rVert}_{w_{x}}+\lVert{\mathbf{u}[i]\rVert}_{w_{u}}+J_{o}(\mathbf{x}[i],\mathbf{p}_{o}[i])),$ (1a)
@@
$\frac{dx}{dt}=v\cos\theta,\frac{dy}{dt}=v\sin\theta,\frac{dv}{dt}=a,\frac{d\theta}{dt}=\omega.$ (1b)
Here $\mathbf{x}_{r}$ is a reference path, $\mathbf{p}_{o}$ is a vector of obstacle parameters. The total cost function consist of three terms: path following term ($\lVert{\mathbf{x}[i]-\mathbf{x}_{r}[i]\rVert}_{w_{x}}$ is a weighted distance between actual trajectory and a reference path), control minimization term ($\lVert{\mathbf{u}[i]\rVert}_{w_{u}}$ is a weigted norm of the control input), and repulsive potential $J_{o}$. Aforementioned statement may be adopted to systems with other dynamic models (e.g., holonomic or car-like robots): this requires the change of $\mathbf{x}$, $\mathbf{u}$, and (1b), while the obstacle model will keep the same. Neural potential function is a neural network, which is trained to predict $J_{o}$ based on $\mathbf{x}$ and $\mathbf{p}_{o}$. In NPField $\mathbf{p}_{o}$ is an embedding of the obstacle map obtained from the neural encoder. Reference potential for each point is calculated based on signed distance ($SDF$) to obstacle border: $J_{o}=w_{1}(\pi/2+arctan(w_{2}-w_{2}SDF))$ ($w_{1}$ and $w_{2}$ are weighting coefficients). The reference potential for the whole robot is chosen as the maximum potential of the points within its footprint. Reference potential is non-differentiable as $SDF$, and maximum values are calculated algorithmically. Therefore, it cannot be directly used in the MPC loop. Instead, reference potential is used to generate a dataset for training the neural potential field. The difference between static and dynamic environments is that in the first case $\mathbf{p}_{o}$ is constant for the whole trajectory, while in the second case it depends on time.

### v3

@@ Abstract
Generalist robot policies must operate safely and reliably in everyday human environments such as homes, offices, and warehouses, where people and objects move unpredictably. We present Dynamic Neural Potential Field (NPField-GPT), a learning-enhanced model predictive control (MPC) framework that couples classical optimization with a Transformer-based predictor of footprint-aware repulsive potentials. Given an occupancy sub-map, robot footprint, and optional dynamic-obstacle cues, our NPField-GPT model forecasts a horizon of differentiable potentials that are injected into a sequential quadratic MPC program via L4CasADi, yielding real-time, constraint-aware trajectory optimization. We additionally study two baselines: NPField-StaticMLP, where a dynamic scene is treated as a sequence of static maps; and NPField-DynamicMLP, which predicts the future potential sequence in parallel with an MLP.
@@
In dynamic indoor scenarios from BenchMR and on a Husky UGV in office corridors, NPField-GPT produces more efficient and safer trajectories under motion changes, while StaticMLP/DynamicMLP offer lower latency. We also compare with the CIAO* and MPPI baselines. Across methods, the Transformer+MPC synergy preserves the transparency and stability of model-based planning while learning only the part that benefits from data: spatiotemporal collision risk. Code and trained models are available at https://github.com/CognitiveAISystems/Dynamic-Neural-Potential-Field.
I Introduction
@@
Mobile robots that act autonomously within human-oriented environments can become significant assistants for humans. They can be applied in offices, shops, homes, and medical facilities. Autonomous operation requires methods for planning motion within the environment. Trajectory planning is often executed in two stages: first, a rough global path is generated via search-based [9, 25] or sampling-based [16, 20] methods; second, the global path is turned into a local trajectory under kinodynamic constraints and obstacle avoidance. The second stage is often executed online with a receding horizon strategy. Solutions for receding-horizon trajectory planning may be obtained with model predictive path integral (MPPI) [39, 40] or numerical model predictive control (MPC) [5, 14, 33, 34, 44, 6, 42, 37]. MPPI can work with arbitrary obstacle maps, but it can produce invalid or chattering solutions. Numerical MPC provides fast and stable solutions; however, it requires an analytical representation of collision danger either as a set of constraints [33] or as a cost term, the repulsive potential. The term “repulsive potential” was introduced for the artificial potential field global planner [17]. In MPC, the trajectory is biased toward safer regions according to the repulsive potential. In this work, we consider local planning with numerical MPC in the presence of dynamic obstacles (Fig. 1).
@@
Figure 1: Overview of NPField-GPT integrated with MPC for dynamic obstacle avoidance. Inputs include occupancy map, robot footprint, robot pose, and dynamic-obstacle state $\left(x_{dyn},y_{dyn},\theta_{dyn}\right)$. The neural model predicts a footprint-aware repulsive potential horizon, which is consumed by MPC to optimize $\mathbf{x}_{opt}$ and $\mathbf{u}_{opt}$ in real time.
@@
Avoiding collisions with dynamic obstacles (e.g., walking human beings in an office environment) is especially challenging. It means that the obstacle map for each frame within the prediction horizon will be different. The prediction of obstacle flow is a distinct task [13, 35, 24], which is out of our scope. It can be solved with various methods. Here, we consider dynamic obstacles to be predictable. The dynamic appearance of the obstacle model enlarges the computational complexity of the MPC problem due to a higher number of obstacle parameters.
We follow a hybrid design principle: learn only spatiotemporal collision risk and keep optimization model-based. Specifically, we extend NPField to dynamic scenes with three neural variants that produce differentiable, footprint-aware obstacle potentials for MPC.
Our framework comprises three architectures to balance accuracy and latency. The simplest variant, StaticMLP, processes dynamic environments as sequential static frames. To improve temporal reasoning, DynamicMLP infers future potentials concurrently using parallel heads. Finally, NPField-GPT employs a non-autoregressive Transformer to predict the entire potential horizon holistically, significantly enhancing the representation of spatial and temporal obstacle dynamics. In the GPT model, coordinate information is emphasized through explicit dynamic-obstacle channels, relative-coordinate fusion, and near-obstacle weighted training.
Our main contributions are:
  • A real-time MPC-compatible dynamic neural potential framework with three variants (StaticMLP, DynamicMLP, GPT) covering accuracy-latency trade-offs.
  • A novel NPField-GPT architecture that predicts horizon potentials in parallel and strengthens obstacle-relative geometry conditioning.
  • Extensive comparisons with CIAO* and MPPI in BenchMR scenarios and real Husky UGV office-corridor experiments.
  • Open-source implementation with L4CasADi integration for differentiable neural costs inside MPC.
II Related works
@@
II-A Receding horizon planning and dynamic obstacles
@@
The task of local planning is to turn a fixed part of the rough global plan into a smooth trajectory segment under consideration of obstacles and kinodynamic constraints. Model Predictive Path Integral (MPPI) [39] achieves this via sampling random trajectory segments and generating a good solution based on these samples. Collision check for sample trajectories may be systematically extended to the case of dynamic obstacles [23, 29]. Computational heaviness and nondeterministic nature are disadvantages of MPPI; therefore, it may be used in cases when the process model is too complicated for MPC. Numerical MPC solves the local planning task as an optimal control problem. Obstacle avoidance is formalized as a set of constraints (e.g. [33]) or as an additional cost term of the optimization problem. This formalization leads to additional problem parameters describing obstacle properties. The number of parameters should be small to preserve MPC performance. Many approaches approximate the obstacle map with a low-dimensional model. Either obstacles or free space may be approximated with a set of simple geometric figures: points [14], circles [33, 42], rectangles [34], polylines [43], or polygons [5, 8, 37]. Some of these works explicitly consider dynamic environments [43, 8]; others can be adapted straightforwardly (e.g., CIAO* [33, 34] uses independent approximations of the free space at each timestep; these approximations can reflect dynamic obstacles). The limitations of geometric approximations are computational challenges and loss of fidelity. Many works (e.g. [5, 37]) do not consider how to obtain a geometric approximation from an arbitrary obstacle map. In practice, maps are often provided as occupancy grids (a matrix projected onto the map; zero values correspond to free-space cells, while ones correspond to occupied cells). High-resolution grids have too many parameters to be passed to the MPC solver in raw form. In the next subsection, we consider approaches where obstacle data are modeled with neural networks.
II-B Trajectory optimization with neural collision model
@@
Learning collision model for numeric trajectory optimization was considered in several works [1, 31, 18, 15, 12]. It is challenging to balance the precision and computational complexity of the collision model; therefore, existing works are constrained in terms of computation time and complexity of the maps. [1, 31, 18, 15] use neural models inspired by Neural Radiance Field [22]. These models learn the structure of a single obstacle map (in 2D or 3D) and may be used for navigation within the learned map. [1, 31] exploit network inference for trajectory optimization, while [18, 15] optimize trajectory within the learning procedure. Instead of learning a single obstacle map, the models from [2, 12] take an obstacle representation as input. Both architectures include two submodels: the first reduces the dimensionality of the obstacle representation, while the second calculates the collision score. Therefore, the first submodel provides a compact vector of parameters for a real-time MPC solver, while the second is integrated into this solver using the L4CasADi [31] framework. The difference is that NPField [2] utilizes occupancy grids for footprint-aware collision avoidance of a wheeled mobile robot, while [12] utilizes depth images for 3D collision avoidance of an aerial robot. To our knowledge, there are no neural obstacle models that provide MPC avoidance from dynamic obstacles in real time. [15] exploits a NeRF learning procedure for trajectory optimization in the presence of moving obstacles; however, the computational procedure takes tens of seconds, which is non-real-time. This work aims to provide a model that allows online MPC collision avoidance with a neural model of static and dynamic obstacles.
II-C Positioning and novelty.
Unlike classical geometric MPC [33, 34] or sampling-based methods [39], our approach naturally encodes footprint-aware costs from raw occupancy grids. Furthermore, in contrast to prior neural models that learn single scenes offline [1, 18, 15] or use non-map inputs [12], our framework provides real-time, differentiable repulsive potentials for dynamic obstacle avoidance directly within the MPC loop.
III Background
@@
Following NPField, we define the statement and notations for model predictive local planning. We consider a non-holonomic wheeled robot with a differential drive. System state vector $\mathbf{x}=\{x,y,\theta,v\}$ includes 2D robot coordinates $x$ and $y$, its orientation $\theta$ and linear velocity $v$ (directed according to $\theta$). Control vector $\mathbf{u}=\{a,\omega\}$ includes robot acceleration $a$ (directed according to $\theta$) and angular velocity $\omega$. We consider the model with continuous time dynamics and discrete time control (i.e., $u$ is considered to be constant within the single timestep). A formal statement for the trajectory optimization problem is the following:
$\{\mathbf{x}_{opt}[i],\mathbf{u}_{opt}[i]\}_{i=k}^{k+m}=$ (1a)
$\arg\min\sum_{i=k}^{k+m}(\lVert{\mathbf{x}[i]-\mathbf{x}_{r}[i]\rVert}_{w_{x}}+\lVert{\mathbf{u}[i]\rVert}_{w_{u}}+J_{o}(\mathbf{x}[i],\mathbf{p}_{o}[i])),$
@@
$\frac{dx}{dt}=v\cos\theta,\quad\frac{dy}{dt}=v\sin\theta,\quad\frac{dv}{dt}=a,\quad\frac{d\theta}{dt}=\omega.$ (1b)
Here $\mathbf{x}_{r}$ is a reference path, $\mathbf{p}_{o}$ is a vector of obstacle parameters. The total cost function consists of three terms: path following term ($\lVert{\mathbf{x}[i]-\mathbf{x}_{r}[i]\rVert}_{w_{x}}$ is a weighted distance between actual trajectory and a reference path), control minimization term ($\lVert{\mathbf{u}[i]\rVert}_{w_{u}}$ is a weighted norm of the control input), and repulsive potential $J_{o}$. The aforementioned statement may be adapted to systems with other dynamic models (e.g., holonomic or car-like robots): this requires the change of $\mathbf{x}$, $\mathbf{u}$, and (1b), while the obstacle model will keep the same. Neural potential function is a neural network, which is trained to predict $J_{o}$ based on $\mathbf{x}$ and $\mathbf{p}_{o}$. In NPField $\mathbf{p}_{o}$ is an embedding of the obstacle map obtained from the neural encoder. Reference potential for each point is calculated based on the signed distance function (SDF) to the obstacle border. The reference potential for the whole robot is chosen as the maximum potential of the points within its footprint. The reference potential is non-differentiable (as is the SDF), and the maximum over footprint cells is computed algorithmically, so it cannot be used directly in the MPC loop. Instead, reference potential is used to generate a dataset for training the neural potential field. The difference between static and dynamic environments is that in the first case $\mathbf{p}_{o}$ is constant for the whole trajectory, while in the second case it depends on time.
Figure 2: StaticMLP (left), DynamicMLP (middle), and NPField-GPT (right). Yellow: map/footprint encoder; green: potential predictor for sampled states; red: auxiliary decoder used only during training. NPField-GPT predicts horizon potentials in one Transformer pass using learnable query tokens and coordinate-enhanced conditioning.